OpenStudy (anonymous):

for two independent events events A and B , P(A)=0.40 P(B)=0.35 Find P(A and B)

OpenStudy (nirmalnema):

p(A and B)=pA*pB =0.4*0.35=0.14

OpenStudy (anonymous):

thanks @nirmalnema

OpenStudy (nirmalnema):

your welcome..!!